We have total 30 balls which are grouped by size and color as follows. What is the probability that a randomly chosen ball is not large or not yellow? Large Small Red 5 5 Blue 4 6 Yellow 7 3
Answer #1
Large Small
Red 5 5
Blue 4 6
Yellow 7 3

P(a randomly chosen ball is not large or not yellow) = 1 - P(a randomly chosen ball is large and yellow)

= 1 - 7/30

= 23/30